# Are all models wrong?Fundamental limits in distribution-free empirical model falsification

### v1

@@ B.1 Proof of Theorem 6 (piecewise constant functions)
Before presenting the proof, we first describe the intuition at a high level. Consider any $f\in\mathcal{F}_{\mathrm{pwc}}^{(m)}$, which takes (at most) $m$ distinct values in its output. If $m$ is not too large, then with high probability, at most $n-1$ of these values will actually be observed in a random sample of size $n$ (that is, sampling $n$ covariates $X_{1},\ldots,X_{n}$ will likely result in at most $n-1$ distinct values amongst $f(X_{1}),\ldots,f(X_{n})$). Consequently, on the sample $\mathcal{D}_{n}$, the empirical risk of $f$ can also be achieved by some function in $\mathcal{F}_{\mathrm{pwc}}^{(n-1)}$, which is a much less complex model class. As an interesting note, we remark that the lower bound can be expressed as
$\hat{L}_{\alpha}^{\mathrm{pwc}}(\mathcal{F}_{\mathrm{pwc}}^{(m)},\mathcal{D}_{n})=\hat{L}_{\alpha_{1}}^{\mathrm{ERM}}(\mathcal{F}_{\mathrm{pwc}}^{(n-1)},\mathcal{D}_{n}),$
that is, the same lower bound as constructed in Theorem 2, but computed with a much lower complexity model class (if $n-1\ll m$).
We are now ready to formally prove the theorem.
Proof of Theorem 6.
@@
Since $\varepsilon>0$ can be taken to be arbitrarily small, this completes the proof. ∎
@@
Since $I$ takes values in the set $[n]$, we have $\mathbb{P}_{P_{Y}}\{I\leq n-1\}=1-\mathbb{P}_{P_{Y}}\{I=n\}$. In order to have $I=n$ we must have all observations $Y_{1},\dots,Y_{n}$ distinct, meaning that
@@
B.2 Extension for the piecewise constant example
Next, we state and prove an extension of Theorem 6, to give a more precise characterization of how the effective complexity $m$ of the model class $\mathcal{F}_{\mathrm{pwc}}^{(m)}$ influences the behavior of the lower bound.
@@ B.2 Extension for the piecewise constant example
$r:=\left\lfloor\left(\frac{n(n-1)}{n+2m}-2\sqrt{\frac{n(n-1)}{n+2m}\log(1/\alpha_{0})}\right)_{+}\right\rfloor.$
@@
denote the number of unique values observed in the sample. Then
@@
$r:=\left\lfloor\left(\frac{n(n-1)}{n+2m}-2\sqrt{\frac{n(n-1)}{n+2m}\log(1/\alpha_{0})}\right)_{+}\right\rfloor.$
@@
$r(Y)=\sum_{j=1}^{m}(C_{j}(Y)-1)_{+}\textnormal{ where }C_{j}(Y)=\sum_{i=1}^{n}\mathbbm{1}_{Y_{i}=y_{j}}.$
Fixing any $k\geq 2$, and defining $Y_{-i}=(Y_{1},\dots,Y_{i-1},Y_{i+1},\dots,Y_{k})\in\{y_{1},\dots,y_{m}\}^{k-1}$, note that $C_{j}(Y)=C_{j}(Y_{-i})$ for all $i$ with $Y_{i}\neq y_{j}$. Therefore, for any $i,j$ with $Y_{i}=y_{j}$,
@@
Next we calculate $\mathbb{E}_{P_{Y}}[r(Y)]$. First, $C_{j}\sim\mathrm{Binom}(n,p_{j})$, where $p_{j}=\mathbb{P}_{P_{Y}}\{Y=y_{j}\}$. Thus
$\mathbb{E}_{P_{Y}}[(C_{j}-1)_{+}]=\mathbb{E}_{P_{Y}}[C_{j}-1+\mathbbm{1}_{C_{j}=0}]=np_{j}-1+(1-p_{j})^{n}.$
@@
Now let $p\in(p_{1},\dots,p_{m})$, which lies in the probability simplex. Since $p\mapsto\sum_{j=1}^{m}(1-p_{j})^{n}$ is convex, this means that $\sum_{j=1}^{m}(1-p_{j})^{n}$ is minimized at $p=(\frac{1}{m},\dots,\frac{1}{m})$, i.e.,
@@
since $t\mapsto(t-2\sqrt{t\log(1/\alpha_{0})})_{+}$ is a nondecreasing function, and $r(Y)$ is nonnegative.
@@ C.1 Proofs for results on the linear model example
First, we define two continuous distributions on $(\mathbf{X},\beta)\in\mathbb{R}^{d\times n}\times\mathbb{R}^{d}$. Let $f(x,y)$ denote the density of the distribution $P$ on $\mathbb{R}^{d}\times\mathbb{R}$. Now, fix any constant $c>0$, and define $\tilde{P}$ as the distribution with density
@@
where $\mathbf{P}_{\mathbf{x}}=\mathbf{x}^{\top}(\mathbf{x}\mathbf{x}^{\top})^{-1}\mathbf{x}\in\mathbb{R}^{n\times n}$ is the projection matrix to the row span of $\mathbf{x}$. (Note that we are implicitly assuming $\mathbf{x}\mathbf{x}^{\top}\in\mathbb{R}^{n\times n}$ is invertible—since we are defining a density, it is sufficient that this holds for almost every $\mathbf{x}\in\mathbb{R}^{n\times d}$.) Next, define $\tilde{Q}$ as the distribution with density
@@
by the definition of $h(\cdot)$. (We will verify below that these functions are indeed well-defined densities, i.e., that $C$ is finite and positive, and that $g_{\tilde{P}}$ and $g_{\tilde{Q}}$ each integrate to $1$.)
Next, let $\tilde{P}_{(\mathbf{X},\mathbf{Y})}$ denote the distribution of $(\mathbf{X},\mathbf{Y}):=(\mathbf{X},\mathbf{X}\beta)$ induced by drawing $(\mathbf{X},\beta)\sim\tilde{P}$, and define $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ analogously. Below, we will prove that
$\tilde{P}_{(\mathbf{X},\mathbf{Y})}=P^{n}$, and $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}\in\tilde{\mathcal{Q}}_{\mathrm{lin}}^{(n)}$. (C.2)
Assuming (C.2) holds, we then have
$\lambda_{n,d}(P)=\inf_{Q\in\tilde{\mathcal{Q}}_{\mathrm{lin}}^{(n)}}\mathrm{d}_{\mathrm{TV}}(P^{n},Q)\leq\mathrm{d}_{\mathrm{TV}}(P^{n},\tilde{Q}_{(\mathbf{X},\mathbf{Y})})\\ =\mathrm{d}_{\mathrm{TV}}(\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})})\\ \leq\mathrm{d}_{\mathrm{TV}}(\tilde{P},\tilde{Q}),$
where the last step holds by construction of $\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ from $\tilde{P},\tilde{Q}$. Moreover,
@@
by definition of $\mathbf{Y}=\mathbf{X}\beta$. From (C.2), we know that $P^{n}=\tilde{P}_{(\mathbf{X},\mathbf{Y})}$, and therefore,
$\lambda_{n,d}(P)\leq\mathrm{d}_{\mathrm{TV}}(\tilde{P},\tilde{Q})\\ \leq\frac{1}{2}\mathbb{E}_{P}\left[\left|\frac{h(\mathbf{X})}{\mathbb{E}_{P}[h(\mathbf{X})]}-1\right|\right]+\mathbb{E}_{P}\left[1-e^{-\|\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{Y}\|^{2}_{2}/(2c)}\right].$
Moreover, since this is true for any $c>0$, and since for any random variable $A\geq 0$ it holds that $\lim_{c\to\infty}\mathbb{E}[e^{-A/c}]=1$ by the dominated convergence theorem, we therefore have
as desired.
To complete the proof, we now need to verify (C.2) (and, along the way, to check that $g_{\tilde{P}}$ and $g_{\tilde{Q}}$ are well-defined densities). We begin by considering a joint distribution $\tilde{P}_{*}$ on $(\mathbf{X},\mathbf{Y},\beta)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}\times\mathbb{R}^{d}$ generated as follows: first sample $(\mathbf{X},\mathbf{Y})\sim P^{n}$, then sample
@@
which means that $g_{\tilde{Q}}$ must also integrate to $1$ (i.e., it is a well-defined density). Next we need to verify that $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}\in\tilde{\mathcal{Q}}_{\mathrm{lin}}^{(n)}$. Define $\tilde{Q}_{\beta}$ as the distribution of $(\mathbf{X},\mathbf{Y})$ conditional on $\beta$, under the joint distribution $\tilde{Q}$ on $(\mathbf{X},\beta)$, when we define $\mathbf{Y}=\mathbf{X}\beta$. Clearly, $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ can be expressed as a mixture of such distributions. Moreover, the data points $(X_{i},Y_{i})$ are i.i.d. conditional on $\beta$ (since by examining the density $g_{\tilde{Q}}$ we can see that it factors over data points $i$), and satisfy $Y_{i}=X_{i}^{\top}\beta$ almost surely. Therefore, $\tilde{Q}_{\beta}\in\mathcal{Q}_{\mathrm{lin}}^{(n)}$, which completes the proof. ∎
@@
where the second step holds by Pinsker’s inequality (to be more concrete, we are using the fact that, for a random variable $Z\geq 0$ with $\mathbb{E}[Z]=1$, $\frac{1}{2}\mathbb{E}[|Z-1|]\leq\sqrt{\frac{1}{2}\log(1/Z)}$). Next we compute each of these expected values. First, observe that $\mathbf{X}\Sigma^{-1}\mathbf{X}^{\top}\sim W_{n}(\mathbf{I}_{n},d)$ (a Wishart distribution). Goodman, (1963) proves that the determinant of a random matrix drawn from the Wishart distribution $W_{n}(\mathbf{I}_{n},d)$ is distributed as the product of independent random variables with a $\chi^{2}$-distribution, i.e.,
@@
C.2 Extension for the linear model example
In this section, we extend the results of Section 4.2, to prove a bound on $\lambda_{n,d}(P)$ for a broader family of distributions.

### v2

@@ B.1 Proof of Theorem 6 (piecewise constant functions)
Since $\varepsilon>0$ can be taken to be arbitrarily small, this completes the proof of the theorem.
@@
If $m<n$, the result is immediate. Suppose from this point on then that $m\geq n$. Since $I$ takes values in the set $[n]$, we have $\mathbb{P}_{P_{Y}}\{I\leq n-1\}=1-\mathbb{P}_{P_{Y}}\{I=n\}$. In order to have $I=n$ we must have all observations $Y_{1},\dots,Y_{n}$ distinct, meaning that
@@
B.2 Extensions for the piecewise constant example
Next, we state and prove several extensions of Theorem 6, to give a more precise characterization of how the effective complexity $m$ of the model class $\mathcal{F}_{\mathrm{pwc}}^{(m)}$ influences the behavior of the lower bound.
@@ B.2 Extensions for the piecewise constant example
$r:=\left\lceil\left(\frac{n(n-1)}{n+2m}-2\sqrt{\frac{n(n-1)}{n+2m}\log(1/\alpha_{0})}\right)_{+}\right\rceil.$
@@
denote the number of unique values observed in the sample. Then for any $\alpha_{0}\in(0,1)$,
@@
$r:=\left\lceil\left(\frac{n(n-1)}{n+2m}-2\sqrt{\frac{n(n-1)}{n+2m}\log(1/\alpha_{0})}\right)_{+}\right\rceil.$
@@
$r(Y)=\sum_{j=1}^{m}(C_{j}(Y)-1)_{+},$
where $C_{j}(Y)=\sum_{i=1}^{n}\mathbbm{1}_{Y_{i}=y_{j}}$ counts the number of times that the value $y_{j}$ was observed in the vector $Y$. Fixing any $k\geq 2$, and defining $Y_{-i}=(Y_{1},\dots,Y_{i-1},Y_{i+1},\dots,Y_{k})\in\{y_{1},\dots,y_{m}\}^{k-1}$, note that $C_{j}(Y)=C_{j}(Y_{-i})$ for all $i$ with $Y_{i}\neq y_{j}$. Therefore, for any $i,j$ with $Y_{i}=y_{j}$,
@@
Next we calculate $\mathbb{E}_{P_{Y}}[r(Y)]$. First, $C_{j}\sim\mathrm{Binom}(n,p_{j})$, where $p_{j}=\mathbb{P}_{P_{Y}}\{Y_{1}=y_{j}\}$. Thus
$\mathbb{E}_{P_{Y}}[(C_{j}-1)_{+}]=\mathbb{E}_{P_{Y}}[C_{j}-1+\mathbbm{1}_{C_{j}=0}]=np_{j}-1+(1-p_{j})^{n},$
@@
Now let $p:=(p_{1},\dots,p_{m})$, which lies in the probability simplex. Since $p\mapsto\sum_{j=1}^{m}(1-p_{j})^{n}$ is convex, this means that $\sum_{j=1}^{m}(1-p_{j})^{n}$ is minimized at $p=(\frac{1}{m},\dots,\frac{1}{m})$, i.e.,
@@
since $t\mapsto(t-2\sqrt{t\log(1/\alpha_{0})})_{+}$ is a nondecreasing function on $t\geq 0$, and $r(Y)$ is nonnegative. By definition of $r$, and using the fact that $r(Y)$ is integer-valued, we therefore have $\mathbb{P}_{P_{Y}}\{r(Y)\geq r\}\geq 1-\alpha_{0}$.
@@
Theorem B.2 offers a more informative lower bound than our original construction, in Theorem 6, both of which are valid without placing a bound on the loss. However, by combining the proof ideas from Theorem A.2 (which uses a truncated loss) and Theorem B.2, we can obtain the following potentially even more informative lower bound.
Theorem B.4.
Fix $\alpha\in(0,1)$, $n\geq 1$, and $m\geq 1$. Let $\alpha_{0}+\alpha_{1}=\alpha$, with $\alpha_{0},\alpha_{1}>0$. Fix any $B>0$, and define $\Delta_{n,B}\in(0,1]$ as the unique solution of
$-\Delta_{n,B}-\log(1-\Delta_{n,B})=\frac{B\log(1/\alpha_{1})}{n\hat{R}(\mathcal{F}_{\mathrm{pwc}}^{(n-r)},\mathcal{D}_{n};B)},$
where $r$ is defined as in Theorem B.2, and where the truncated empirical risk $\hat{R}(\cdot,\mathcal{D}_{n};B)$ is defined as in Theorem A.2. Then
$\hat{L}^{\mathrm{pwc-trunc},r,B}_{\alpha}(\mathcal{F}_{\mathrm{pwc}}^{(m)},\mathcal{D}_{n}):=(1-\Delta_{n,B})\hat{R}(\mathcal{F}_{\mathrm{pwc}}^{(n-r)},\mathcal{D}_{n};B)$
Proof of Theorem B.4.
By the same argument and using the same notation as in the proofs of Theorems 6 and B.2, we have $\mathbb{P}_{P}\{|I(\mathcal{D}_{n})|\leq n-r\}\geq 1-\alpha_{0}$. Let $\Delta^{\prime}_{n,B}$ be the unique solution to
$-\Delta^{\prime}_{n,B}-\log(1-\Delta^{\prime}_{n,B})=\frac{B\log(1/\alpha_{1})}{n\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)}.$
Therefore, on the event $\{\hat{f}_{\varepsilon}\in\mathcal{F}_{\mathrm{pwc}}^{(n-r)}\}\supseteq\{|I(\mathcal{D}_{n})|\leq n-r\}$, we have $\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)=\hat{R}(\hat{f}_{\varepsilon},\mathcal{D}_{n};B)\geq\hat{R}(\mathcal{F}_{\mathrm{pwc}}^{(n-r)},\mathcal{D}_{n};B)$ as in the proof of Theorem B.2, and therefore
$-\Delta_{n,B}-\log(1-\Delta_{n,B})=\frac{B\log(1/\alpha_{1})}{n\hat{R}(\mathcal{F}_{\mathrm{pwc}}^{(n-r)},\mathcal{D}_{n};B)}\geq\frac{B\log(1/\alpha_{1})}{n\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)}=-\Delta^{\prime}_{n,B}-\log(1-\Delta^{\prime}_{n,B}).$
Thus, on the event that $|I(\mathcal{D}_{n})|\leq n-r$, we have $\Delta^{\prime}_{n,B}\leq\Delta_{n,B}$ and so
$\hat{L}^{\mathrm{pwc-trunc},r,B}_{\alpha}(\mathcal{F}_{\mathrm{pwc}}^{(m)},\mathcal{D}_{n})\leq(1-\Delta_{n,B})\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)\leq(1-\Delta^{\prime}_{n,B})\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B).$
By Lemma A.3, $\mathbb{P}_{P}\big{\{}(1-\Delta_{n,B}^{\prime})\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\big{\}}\geq 1-\alpha_{1}$, and so
$\mathbb{P}_{P}\left\{\hat{L}_{\alpha}^{\mathrm{pwc-trunc},r,B}(\mathcal{F}_{\mathrm{pwc}}^{(m)},\mathcal{D}_{n})\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\right\}\\ \geq\mathbb{P}_{P}\big{\{}(1-\Delta_{n,B}^{\prime})\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\textnormal{ and }|I(\mathcal{D}_{n})|\leq n-r\big{\}}\geq 1-\alpha.$
As in the proof of Theorem A.2, we have $\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\leq R_{P}(\mathcal{F}_{\mathrm{pwc}}^{(m)})+\varepsilon$, so it therefore holds that $\mathbb{P}_{P}\left\{\hat{L}_{\alpha}^{\mathrm{pwc-trunc},r,B}(\mathcal{F}_{\mathrm{pwc}}^{(m)},\mathcal{D}_{n})\leq R_{P}(\mathcal{F}_{\mathrm{pwc}}^{(m)})+\varepsilon\right\}\geq 1-\alpha$, and since $\varepsilon>0$ can be taken to be arbitrarily small, this completes the proof. ∎
@@ C.1 Proofs for results on the linear model example
First, we define two continuous distributions on $(\mathbf{X},\beta)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{d}$. Let $f(x,y)$ denote the density of the distribution $P$ on $\mathbb{R}^{d}\times\mathbb{R}$. Now, fix any constant $c>0$, and define $\tilde{P}$ as the distribution with density
@@
where $\mathbf{P}_{\mathbf{x}}=\mathbf{x}^{\top}(\mathbf{x}\mathbf{x}^{\top})^{-1}\mathbf{x}\in\mathbb{R}^{d\times d}$ is the projection matrix to the row span of $\mathbf{x}$. (Note that we are implicitly assuming $\mathbf{x}\mathbf{x}^{\top}\in\mathbb{R}^{n\times n}$ is invertible—since we are defining a density, it is sufficient that this holds for almost every $\mathbf{x}\in\mathbb{R}^{n\times d}$.) Next, define $\tilde{Q}$ as the distribution with density
@@
where the inequality holds by the definition of $h(\cdot)$. (We will verify below that these functions are indeed well-defined densities, i.e., that $C$ is finite and positive, and that $g_{\tilde{P}}$ and $g_{\tilde{Q}}$ each integrate to $1$.) Next, let $\tilde{P}_{(\mathbf{X},\mathbf{Y})}$ denote the distribution of $(\mathbf{X},\mathbf{Y}):=(\mathbf{X},\mathbf{X}\beta)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ induced by drawing $(\mathbf{X},\beta)\sim\tilde{P}$, and define $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ analogously.
The intuition of the remainder of the proof is the following. First, in Step 1, we will verify that $\tilde{P}$ and $\tilde{Q}$ are close in total variation distance—that is, we will bound $\mathrm{d}_{\mathrm{TV}}(\tilde{P},\tilde{Q})$, which therefore induces a bound on $\mathrm{d}_{\mathrm{TV}}(\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})})$, by construction of $\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ from $\tilde{P},\tilde{Q}$. Next, in Step 2, we will verify that
[…]
and consequently, this will mean that $\lambda_{n,d}(P)\leq\mathrm{d}_{\mathrm{TV}}(\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})})$. Finally, in Step 3, we will take a limit as $c\to\infty$ to complete the proof.
Step 1: bounding the total variation distance.
We calculate
@@
by definition of $\mathbf{Y}=\mathbf{X}\beta$. From (C.2) (which we will verify below), we know that $P^{n}=\tilde{P}_{(\mathbf{X},\mathbf{Y})}$, and therefore,
$\mathrm{d}_{\mathrm{TV}}(\tilde{P},\tilde{Q})\leq\frac{1}{2}\mathbb{E}_{P}\left[\left|\frac{h(\mathbf{X})}{\mathbb{E}_{P}[h(\mathbf{X})]}-1\right|\right]+\mathbb{E}_{P}\left[1-e^{-\|\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{Y}\|^{2}_{2}/(2c)}\right].$ (C.3)
Step 2: proving (C.2).
We now need to verify (C.2) (and, along the way, to check that $g_{\tilde{P}}$ and $g_{\tilde{Q}}$ are well-defined densities). We begin by considering a joint distribution $\tilde{P}_{*}$ on $(\mathbf{X},\mathbf{Y},\beta)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}\times\mathbb{R}^{d}$ generated as follows: first sample $(\mathbf{X},\mathbf{Y})\sim P^{n}$, then sample
@@
which means that $g_{\tilde{Q}}$ must also integrate to $1$ (i.e., it is a well-defined density). Next we need to verify that $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}\in\tilde{\mathcal{Q}}_{\mathrm{lin}}^{(n)}$. Define $\tilde{Q}_{\beta}$ as the distribution of $(\mathbf{X},\mathbf{Y})$ conditional on $\beta$, under the joint distribution $\tilde{Q}$ on $(\mathbf{X},\beta)$, when we define $\mathbf{Y}=\mathbf{X}\beta$. Clearly, $\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ can be expressed as a mixture of such distributions. Moreover, the data points $(X_{i},Y_{i})$ are i.i.d. conditional on $\beta$ (since by examining the density $g_{\tilde{Q}}$ we can see that it factors over data points $i$), and satisfy $Y_{i}=X_{i}^{\top}\beta$ almost surely. Therefore, $\tilde{Q}_{\beta}\in\mathcal{Q}_{\mathrm{lin}}^{(n)}$, which completes the proof of (C.2).
Step 3: combining everything.
Combining our calculations so far, we have shown that
$\lambda_{n,d}(P)\leq\mathrm{d}_{\mathrm{TV}}(\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})})\leq\mathrm{d}_{\mathrm{TV}}(\tilde{P},\tilde{Q})\\ \leq\frac{1}{2}\mathbb{E}_{P}\left[\left|\frac{h(\mathbf{X})}{\mathbb{E}_{P}[h(\mathbf{X})]}-1\right|\right]+\mathbb{E}_{P}\left[1-e^{-\|\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{Y}\|^{2}_{2}/(2c)}\right],$
where the first step holds by (C.2), the second step holds by construction of $\tilde{P}_{(\mathbf{X},\mathbf{Y})},\tilde{Q}_{(\mathbf{X},\mathbf{Y})}$ from $\tilde{P},\tilde{Q}$, and the third step holds by (C.3). Moreover, since this is true for any $c>0$, and since for any random variable $A\geq 0$ it holds that $\lim_{c\to\infty}\mathbb{E}[e^{-A/c}]=1$ by the dominated convergence theorem, we therefore have
as desired. ∎
@@
where the second step holds by Pinsker’s inequality (to be more concrete, we are using the fact that, for a random variable $Z\geq 0$ with $\mathbb{E}[Z]=1$, $\frac{1}{2}\mathbb{E}[|Z-1|]\leq\sqrt{\frac{1}{2}\mathbb{E}[\log(1/Z)]}$). Next we compute each of these expected values. First, observe that $\mathbf{X}\Sigma^{-1}\mathbf{X}^{\top}\sim W_{n}(\mathbf{I}_{n},d)$ (a Wishart distribution). Goodman, (1963) proves that the determinant of a random matrix drawn from the Wishart distribution $W_{n}(\mathbf{I}_{n},d)$ is distributed as the product of independent random variables with a $\chi^{2}$-distribution, i.e.,
@@
C.2 Extensions for the linear model example: generalizing to other distributions
We extend the results of Section 4.2, to prove a bound on $\lambda_{n,d}(P)$ for a broader family of distributions.